OpenStudy (anonymous):

What is the focus of the parabola given by the equation y2 = -24x? A) (-6, 0) B) (0, 0) C) (0, 0) D) (0, -6)

OpenStudy (anonymous):

(0, 0)

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (muzzack):

B and C?

OpenStudy (vishweshshrimali5):

The focus of a parabola with general equation: \[\large{y^2 = -4ax}\] is (-a,0)

OpenStudy (vishweshshrimali5):

SO it should be A

OpenStudy (muzzack):

https://www.desmos.com/calculator/m1inuwjpgu

OpenStudy (muzzack):

@vishweshshrimali5 if it doesnt have a its obvious your wrong

OpenStudy (vishweshshrimali5):



OpenStudy (vishweshshrimali5):

@Muzzack It has an "a": its actually y^2 = -4*6*x Here, a = 6

OpenStudy (vishweshshrimali5):

She is asking the "focus " not the "vertex".